OpenStudy (anonymous):

multiply:

OpenStudy (anonymous):

2*8 = 16

OpenStudy (anonymous):

\[\sqrt{7} * \sqrt{3}\]

OpenStudy (anonymous):

It won't load correctly.

OpenStudy (anonymous):

sqrt(7) * sqrt(3)

OpenStudy (anonymous):

Okay, so there's this rule that says \(\Large \sqrt{a}*\sqrt{b} = \sqrt{a*b} Basically, you can bring them together.

OpenStudy (anonymous):

\(\Large \sqrt{a}*\sqrt{b} = \sqrt{a*b}\)' Closed tag.

OpenStudy (anonymous):

Well, I didn't know. Lol. so they just multiply together?

OpenStudy (anonymous):

\[\sqrt{7 } \times \sqrt{3} = \sqrt{7\times3} = \sqrt{21}\]

OpenStudy (anonymous):

Yes. Now, you should look to see if it can be simplified from there. Is 21 a perfect square? Are there any perfect square factors?

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

Can \(\sqrt{21} \) be simplified further now?

OpenStudy (anonymous):

Nope =) It's as simple as it can be.

OpenStudy (anonymous):

yeah it can be simplified

OpenStudy (anonymous):

Haha whut

OpenStudy (anonymous):

Satellite got me questionin' my knowledge.

OpenStudy (anonymous):

\[\large \sqrt{21}=\sqrt{7}\times \sqrt{3}\]

OpenStudy (anonymous):

ahahahaha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

You frickin' goofball.

OpenStudy (anonymous):

point being that there is no such mathematical operation as "simplify" you think \(\sqrt{21}\) is "simple" and i think \(\sqrt{7}\times \sqrt{3}\) is simple it is a matter of taste

OpenStudy (anonymous):

Now I'm confused.. lol.

OpenStudy (anonymous):

@Mikeyy1992 ignore me i am making a point, but don't let me confuse you \(\sqrt{21}\) is the correct answre

OpenStudy (anonymous):

I agree with you, Satellite. Rewrite is a better term for it.

OpenStudy (anonymous):

fluttering i think

OpenStudy (anonymous):

Yes. It means fluttering.

OpenStudy (anonymous):

@satellite73... very smart yet quite humorous.

OpenStudy (anonymous):

Haha quite smart. Smart enough that when he said it could be simplified, I was convinced enough to try and find a perfect square factor.

OpenStudy (anonymous):

*doesn't close question, just to watch everyone talk* Lol :P

OpenStudy (anonymous):

well really if you were asked to "simplify" wouldn't \(\sqrt{6400}\) be simpler if first you wrote it as \(\sqrt{64}\times \sqrt{100}\) ?

OpenStudy (anonymous):

Yes, certainly. Kind of.

OpenStudy (anonymous):

good point Satellite. But some teachers, people, students say refer it as "simplify", but idk, simply depends on the person.

OpenStudy (anonymous):

"Simplify" is just a very poorly defined term.

OpenStudy (anonymous):

when i am king of the math teachers i will banish the word "simplify" from their vocabulary is \(\frac{1}{3}\) simpler than \(\frac{4}{12}\)? no not if i want to add \(\frac{1}{3}+\frac{7}{12}\) it isn't

OpenStudy (anonymous):

You aren't king of the math teachers yet?

OpenStudy (anonymous):

i have seen questions that say "simplify" \(2x(x+3)\) when it is clear it means "multiply"

OpenStudy (anonymous):

well your kinda already the "king of math teachers" here on OS lol

OpenStudy (anonymous):

When I succeed you as king of the math teachers, I'll ban teaching the "distance formula." And just make sure people actually understand the Pythagorean theorem and how to use it.

OpenStudy (anonymous):

amen to that as well

OpenStudy (anonymous):

also on my hit list is "cancel"

OpenStudy (anonymous):

reduce to 1?

OpenStudy (anonymous):

sometime the \(x\) cancels in \(x^2+x-x-1\)

OpenStudy (anonymous):

sometimes it cancels in \(\frac{x^3}{x^2}=x\)

OpenStudy (anonymous):

SmoothMath, lol why would you ban teaching the "distance formula", just curious on your thought.

OpenStudy (anonymous):

sometimes cancelling means \(\sqrt{x^2}=x\)

OpenStudy (anonymous):

i have even heard "cancel" as in \(e^{\ln(x)}=x\)

OpenStudy (anonymous):

adds to zero reduce to lowest terms compose a function with its inverse

OpenStudy (anonymous):

if you call the damn thing what it is, there will be infinitely less confusion

OpenStudy (anonymous):

and now i will go take my valium

OpenStudy (anonymous):

 So the way we're told to teach kids to find the distance between two points is: "Use the distance formula." There are a number of things wrong with this: 1) They have to actually MEMORIZE the distance formula. Most of them won't, and even if they do, memorization and regurgitation is the lowest form of learning. 2) It encourages them to follow a formulaic, procedural way of solving the problem instead of actually THINKING. 3) By the time they get to this in Geometry, they already HAVE all the tools they need to solve this problem. In fact, it's ONE tool, and it's the Pythagorean theorem. The distance formula IS the pythagorean theorem, so we're making them learn it AGAIN, except in a more complicated looking form.